OpenStudy (anonymous):

[10.06] What is the product of 3the square root of u. (the square root of 448. + 4the square root of 7.)? Simplify if possible.

OpenStudy (anonymous):

\[3\sqrt(u)(\sqrt(448)+4\sqrt(7)) = 3\sqrt(u)* (4\sqrt(7)+ 8\sqrt(7))\] = \[3\sqrt(u)\sqrt(7)(12) = 36\sqrt(7)\sqrt(u)\]

OpenStudy (anonymous):

That looks kind of ugly, but did you follow?

OpenStudy (anonymous):

i had options for answers lke this 24the square root of 7 times u. + 4the square root of 7. 204the square root of 7 times u. 24uthe square root of 7. 36the square root of 7 times u.

OpenStudy (anonymous):

Is this what the problem looked like? \[3\sqrt(u)* (\sqrt(448)+ 4\sqrt(7))\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok is the last answer \[36\sqrt(7u)\]?

OpenStudy (anonymous):

That is the answer if it looks like that because \[\sqrt(7)\sqrt(u) = \sqrt(7u)\]

OpenStudy (anonymous):

thanks for all ur help

OpenStudy (anonymous):

np